Larry's time to travel 208 miles is 1 hour more than Terrell's time to travel 165 miles. Terrell drove 3 miles per hour faster than Larry. How fast did each one travel?
Answer by stanbon(75887)

Larry's time to travel 208 miles is 1 hour more than Terrell's time to travel 165 miles. Terrell drove 3 miles per hour faster than Larry. How fast did each one travel?
-----
Larry DATA:
distance = 208 miles ; time = x+1 hrs ; rate = d/t = 208/(x+1) mph
------
Terrell DATA:
distance = 165 miles ; time = x hrs ; rate = d/t = 165/x mph
---------
Equation:
Terrell rate - Larry rate = 3 mph
165/x - 208/(x+1) = 3
------
165(x+1) - 208x = 3x(x+1)
-----
165x + 165 - 208x = 3x^2+ 3x
-----
3x^2 + 3x + 43x - 165 = 0
3x^2 + 46 x - 165 = 0
Positive solution:
x = 3
-----
Terrell rate:: 165/3 = 55 mph
Larry rate::::208/(4) = 52 mph
==================================
Cheers,
Stan H.